OpenStudy (clamin):

PLEASE HELP!! i keep figuring this out but im kind of confused 1. Calculate the number that you need to add to each side of the equation w^2 - 3w = 350 to create a perfect square trinomial. 2. Factor the trinomial. 3. Now that you've factored the equation, find the square root of each side and solve for w.

OpenStudy (anonymous):

It's completing the square, is it not?

OpenStudy (anonymous):

For completing the square problems, always start out with generic square and multiplying it:\[ (x+s)^2=x^2+2sx+s^2 \]The term we will add to both sides is \(s^2\). Now if we have an equation: \[ x^2+bx+c \]We can say \(b=2s\) and thus \(s^2 = (b/2)^2\). We add \((b/2)^2\) to both sides.

OpenStudy (anonymous):

You've provided an equation where \(b=-3\), so we need to add \((-3/2)^2=9/4\) to both sides.

OpenStudy (anonymous):

Well, you can just memorize the \((b/2)^2\) term, but multiplying a generic square allows you to derive it rather quickly.

OpenStudy (clamin):

@wio i got this w^2-3w=350 w^2-3w +(-3/2)^2=350 w^2-3w +(9/4) =350 +(9/4)

OpenStudy (anonymous):

Now remember that \(s=-3/2\), so that means:\[ w^2-3w+(9/4) = (w-3/2)^2 \]

OpenStudy (anonymous):

This comes from our equation \[ x^2+2s+s^2=(x+s)^2 \]but we use \(w\) instead of \(x\).

OpenStudy (anonymous):

If we take the square root of both sides, we get: \[ |w-3/2| = \sqrt{350+9/4} \]Which means \[ w = -3/2\pm \sqrt{350+9/4} \]You should simplify the square root though.

OpenStudy (clamin):

dont u have to square the -3/2

OpenStudy (anonymous):

No, why would you need to square it?

OpenStudy (clamin):

@wio i got this w^2 - 3/2 = square root of 352.25

OpenStudy (clamin):

What should i do nxt??